OpenStudy (highschoolmom2010):

find the geometric mean 3,48 7,9 4,49

OpenStudy (amistre64):

multiply them together and take the square root of the result

OpenStudy (highschoolmom2010):

ok so for 3,48 \[3*48=144\] \[\sqrt{144}=12\]

OpenStudy (mathstudent55):

correct

OpenStudy (highschoolmom2010):

7*9=63 \[\sqrt{63}=\] \[3\sqrt{7}\]

OpenStudy (mathstudent55):

correct

OpenStudy (highschoolmom2010):

ok and for the last one 4*49=196 \[\sqrt{196}=14\]

OpenStudy (rsadhvika):

correct ! 4*49 2*7 14

OpenStudy (rsadhvika):

as you do more problems, you will see that square root of 4 = 2 square root of 49 = 7 so geometric mean = 2x7 = 14

OpenStudy (highschoolmom2010):

i know that i just really tbh didnt want to simplify them :DD

OpenStudy (rsadhvika):

cool /

OpenStudy (highschoolmom2010):

ty @rsadhvika & @mathstudent55 & @amistre64 for the help

OpenStudy (rsadhvika):

yw :)